Fly ball. The approximate distance in feet that a baseball travels when hit at an angle of 45° is given by the expression

(v0)^2/32
where v0 is the initial velocity in feet per second. If Barry Bonds of the Giants hits a ball at a 45° angle with an initial velocity of 120 feet per second, then how far will the ball travel?

Answer by fractalier(6550)

Yes, maximum range of a projectile occurs when shot at 45 degrees...its range is given by
R = v^2 / g
Here, then
R = (120)^2 / 32 = 450 feet